破壊 · 構造力学特論 2 単位 (選択)

Advanced Fracture and Structural Mechanics  村上理一 · 教授/ 知的力学システム工学専攻機械創造システム工学コース生産システム講座

米倉大介 · 准教授/ 知的力学システム工学専攻機械創造システム工学コース生産システム講座

【授業目的】機械 · 構造物の設計や安全性あるいは健全性を確保するために材料 ; に内在するき裂状欠陥の先端に生じる応力分布を力学的に記述するパラメー ; ターとき裂進展に対する材料の抵抗値を定量的に表す内容を講義し，演習，レ ; ポートを実施して機械 · 構造物の安全設計や破壊防止に必要な知識を修得さ ; せる． ; 【授業概要】破壊力学の目的は機械や構造物に使用する部材の強度評価を行い， ; 機械や構造物の安全性を保証することである．そのために部材に発生するき ; 裂の力学的挙動，き裂先端応力解析，破壊力学による破壊解析，さらに弾塑性 ; 状態におけるき裂の非線形力学挙動などを中心に講義を進め，機械や構造物 ; の設計あるいは使用中の破壊に対する防止のための理論的な根拠を平易に講 ; 述する．講義項目は次のとおりである．講義項目ごとに演習あるいはレポー ; トを課し，理解を促す．工業にかかわる科目である． ; 【授業形式】 [形態] ; 【キーワード】破壊，安全設計，き裂，機械 · 構造物，応力拡大係数 ; 【先行科目】 [先行科目] ; 【関連科目】『固体力学』 (0.5)，『材料工学』 (0.5)，『材料物性特論』 (0.5) ; 【履修要件】 [要件] ; 【履修上の注意】 [注意] ; 【到達目標】 ; 1. 破壊力学の概念の理解 ; 2. 応力拡大係数の破壊解析への応用 ; 3. 破壊靱性の物理的意味の理解 ; 4. J 積分の定義と物理的意味の理解 ; 【授業計画】 ; 1. 線形破壊力学の基礎と適用限界 ; 2. 応力拡大係数の評価 ; 3. 破壊靱性と破壊解析 ; 4. 破壊力学による機械安全性設計 ; 5. き裂先端付近の弾塑性変形 ; 6. 線形破壊力学と非線型破壊力学 ; 7. 中間試験 ; 8. J 積分の定義と性質 9. J 積分の評価方法

10. J 積分による破壊靱性評価

11. 破壊開始の条件

12. き裂の安定成長と不安定破壊

13. 非線形破壊力学の適用例

14. 非線形破壊力学による破壊解析

15. 破壊力学のまとめ

16. 定期試験

【成績評価基準】グループ討論，レポートおよび試験によって，目標 3 から 6 の項

目について目標達成ができていることを確認し，平均 60%以上が合格【教科書】プリント【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216811 【対象学生】 [対象学生] 【連絡先】

⇒ 村上 (M318, 088-656-7392, murakami@me.tokushima-u.ac.jp) ���� (オフィスアワー: 月曜日 4時から5時まで)

Advanced Fracture and Structural Mechanics 2 units (selection)

Ri-ichi Murakami · Professor / Production Systems Engineering, Mechanical Engineering, Intelligent Structures and Mechanics Systems Engineering Daisuke Yonekura · Associate Professor / Production Systems Engineering, Mechanical Engineering, Intelligent Structures and Mechanics Systems Engineering

Target⟩ Because the machine and structure should be safety, it is important to evaluate the stress distribution in front of the defect like crack which is in the component of machine. In Fracture Mechanics, the stress distribution is expressed by stress intensity factor, K. Also, the fracture of component depends on fracture toughness of material. In this lecture, it is explained that for the defect like crack, how the stress intensity factor is expressed. It is then mentioned that how the fracture is toughness evaluated. The students should submit the reports each exercises. Finaly, the knowledge of safety design and fracture prevention for machine and structure can be mastered.

Outline⟩ The purpose of fracture mechanics is the evaluation of the strength and fracture of machine and structure and then is the guarantee of safety of machine and structure. We should study the mechanical behavior of crack, the stress distribution in front of crack, the fracture analysis by fracture mechanics and the nonlinear mechanical behavior of elastic-plastic state. It is explained the theory of safety design and prevention of fracture for machine and structure.

Style⟩ [形態] �������⟩ fracture, safety design, crack, machine/structure, stress intensity factor Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Solid Mechanics”(0.5), “Material Engineering”(0.5), “Physi-

cal properties of materials”(0.5) Requirement⟩ [要件] Notice⟩ [注意]

Goal⟩

1. Concept of Fracture Mechanics

2. Application to fracture anaysis of stress intensity factor

3. Concept of Fracture Toughness

4. Concept of J Integral

Schedule⟩

1. Introduction of linear fracture mechanics

2. Stress intensity factor

3. Fracture toughness and fracture analysis

4. Safety design of machine by fracture mechanics 5. Elastic-plastic deformation in front of crack tip

6. Linear and nonlinear fracture mechanics

7. Midterm examination

8. Concept of J integral

9. Evaluation method of J integral

10. Fracture toughness by J integral

11. Condition of fracture initiation

12. Stabe and unstable fracture

13. Application of nonlinear fracture mechanics

14. Fracture analysis by nonlinear fracture mechanics

15. Summary

16. Final examination

Evaluation Criteria⟩ Over 60% of group discussion, reports and examination Textbook⟩ Prints Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216811 Student⟩ [対象学生] Contact⟩

⇒ Murakami (M318, +81-88-656-7392, murakami@me.tokushima-u.ac.jp) ���� (Office Hour: Monday 16:00-17:00)